OpenStudy (anonymous):

What is the sum of the first n terms of this series? 4 + 6 + 8 + 10 + 12 + ⋯

geerky42 (geerky42):

Hmm. Would you know the answer if sum is 1+2+3+4+... instead?

OpenStudy (anonymous):

Idk

OpenStudy (zzr0ck3r):

I would do what @geerky42 said \(1+2+3+4+...+n = \dfrac{n(n+1)}{2}\) Multiply every term in that \(\uparrow\) equation by \(4\)

geerky42 (geerky42):

4?

geerky42 (geerky42):

I would say multiply that by 2.

geerky42 (geerky42):

Be careful here. term "2" is missing!

geerky42 (geerky42):

2(1+2+3+4+...) = 2+4+6+8+... But since 2 is missing, you just subtract 2. 2+4+6+8+... - 2 = 4+6+8+...

OpenStudy (anonymous):

Can u explain little bit further I'm still lost

geerky42 (geerky42):

Do you get what @zzr0ck3r said?

OpenStudy (anonymous):

Not really

geerky42 (geerky42):

You don't know the formula 1+2+3+...+n = \(\dfrac{n(n+1)}{2}\)?

OpenStudy (anonymous):

No these problems are ones I have trouble with

geerky42 (geerky42):

oh well, let's just stick to fact that \(1+2+3+4+...+n = \dfrac{n(n+1)}{2}\) For now.

geerky42 (geerky42):

OR do you want me to explain why this is true?

OpenStudy (anonymous):

Yes please

OpenStudy (zzr0ck3r):

Sorry I thought it was 4,8,12,...